OpenStudy (anonymous):

What is the volume of a rectangular crate that has dimensions 9 inches by 9 inches by 1 1/2feet? a.1,458 in.3 b.121.5 in.3 c.1,012.5 in.3 d.36 in.3

OpenStudy (anonymous):

@DaBest21 @welshfella

OpenStudy (anonymous):

@misssunshinexxoxo

OpenStudy (anonymous):

@vbroberson

OpenStudy (anonymous):

@phi

OpenStudy (phi):

first, can you change 1.5 feet into inches?

OpenStudy (anonymous):

no

OpenStudy (phi):

there are 12 inches in 1 foot write the ratio times the number of feet \[ \frac{12 \ inches}{1 \ foot} \cdot 1.5 \ feet\] now multiply can you do that ?

OpenStudy (anonymous):

what numbers to multiply?

OpenStudy (anonymous):

do u know the answer?

OpenStudy (phi):

1 1/2 (or 1.5) can also be written as 3/2 so you can multiply two fractions \[ \frac{12}{1} \cdot \frac{3}{2} \]

OpenStudy (anonymous):

i still can't get the answer

OpenStudy (phi):

are you saying you don't know how to multiply top times top and bottom times bottom in \[ \frac{12}{1} \cdot \frac{3}{2} \]?

OpenStudy (anonymous):

cross multiplication?

OpenStudy (phi):

no, when you multiply two fractions you do top times top bottom times bottom to get a new fraction

OpenStudy (anonymous):

whats the answer the choices are at the top

OpenStudy (phi):

if you would first change 1 1/2 feet into inches, we can get to the problem

OpenStudy (anonymous):

idk how

OpenStudy (anonymous):

I'm just going to choose an answer i think